HETAV PATEL

Grade 11,

ydx+xdy=xy(dy-dx) find differential equation of given question

ydx+xdy=xy(dy-dx) find differential equation of given question

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
hello niraj, the ques is ambiguous as u have asked to find diff eqn which is already given.
so i am assuming you wanted to ask to solve this diff eqn. note the following:
d(xy)= xdy+ydx and d(y – x)=dy – dx 
so ydx+xdy=xy(dy-dx) becomes 
d(xy)/xy= d(y – x)
integrate both sides
ln|xy|= y – x + C, where C is a constant.
or xy= Ke^(y – x)
kindly approve :)